# Collision-inclusive Manipulation Planning for Occluded Object Grasping via Compliant Robot Motions

### v1

@@ Abstract
Traditional robotic manipulation mostly focuses on collision-free tasks. In practice, however, many manipulation tasks (e.g., occluded object grasping) require the robot to intentionally collide with the environment to reach a desired task configuration. By enabling compliant robot motions, collisions between the robot and the environment are allowed and can thus be exploited, but more physical uncertainties are introduced. To address collision-rich problems such as occluded object grasping while handling the involved uncertainties, we propose a collision-inclusive planning framework that can transition the robot to a desired task configuration via roughly modeled collisions absorbed by Cartesian impedance control. By strategically exploiting the environmental constraints and exploring inside a manipulation funnel formed by task repetitions, our framework can effectively reduce physical and perception uncertainties. With real-world evaluations on both single-arm and dual-arm setups, we show that our framework is able to efficiently address various realistic occluded grasping problems where a feasible grasp does not initially exist.
@@ I Introduction
Traditional robotic manipulation research in motion and grasp planning relies strongly on the assumption that a collision-free solution exists. In those paradigms, the robot must compute and then effectively follow a path through free space, avoiding all collisions, in order to reach the desired task configuration [1, 2, 3]. However, in many practical scenarios (e.g., grasping a wide object on a flat surface, and tightly stacked books) showcased in Fig. 1, a collision-free grasp configuration does not initially exist due to occlusions by surrounding objects, and traditional collision-free strategies cannot find a solution. In contrast, like how humans grasp in such configurations, by allowing collisions and safe interactions between the gripper and the environment, the robot can insert its fingers into narrow gaps between objects to create free space for a desired grasp. For example, in Fig. 1, the robot can insert one finger in between the object and the table surface to scoop-grasp a wide object; or insert both fingers into gaps between books to grasp the target one.
@@
Figure 1: Illustrations of occluded grasping tasks solved by our proposed collision-inclusive planner. Top: A book is lying flat on a table surface, needing the robot gripper to scoop-grasp it by contacting the table. Bottom: A target book is tightly surrounded by other books on the shelf, requiring the robot gripper to insert its fingers into the gaps between books for a stable grasp.
@@
Motivated by the insights above, in this work, we propose a framework for occluded object grasping for which the feasible grasp configuration is initially blocked by the environment. Given mesh models of the robot gripper and the objects, our collision-inclusive framework is able to:
@@
    robustly move the robot to the desired task configuration without accurate system modeling and perception;
@@
    generate effective compliant motions that exploit environmental constraints to reduce task uncertainties;
@@ II Related Work
Occluded Object Grasping. When all feasible grasp configurations for the target object are occluded by the environment or surrounding objects, traditional collision-free motion and grasp planning frameworks cannot solve the problem. In such scenarios, the robot needs to reconfigure the objects to certain desired configurations where a grasp can be achieved. To this end, object rearrangement via nonprehensile actions such as pushing can be incorporated for pre-grasp manipulation [10, 11]. By manually designed primitives or learned policies, other motion skills such as sliding the object to the table edge [12, 13] or the wall [14] and flipping the object against the wall [15, 16, 17] can also be leveraged. Unlike the existing strategies which require significant reconfiguration of the objects, this work addresses the occluded grasping problem in a confined space without much object motions, via collision-inclusive compliant actions.
Manipulation Funnels. The idea of manipulation funnels was originally proposed in [9] to robustly manipulate objects despite variations in their initial locations and shapes. The idea was later followed by other early works for peg-in-hole [18], door-closing [19], and parts-feeding [20] applications. However, their funnels were limited to only geometric funnels. More recent work has explored manipulation funnels by empirical analysis [21], and robust in-hand manipulation can be accomplished via the compositions of funnels [22]. In our framework, we broaden the idea of funnels and explore the manipulation funnel formed by repeated actions: With variations in the initial manipulation states (e.g., different gripper poses relative to the object), task repetitions will implicitly eliminate the involved uncertainties and produce stable outcomes (e.g., a desired grasp) after a few trials.
Exploitation of Environmental Constraints. Exploiting environmental constraints (e.g., through motions in contact with the environment), also known as extrinsic dexterity [23], has been extensively investigated to reduce execution uncertainties during manipulation [24]. Such a concept has been successfully applied in various manipulation problems to generate robust solutions in grasping [8, 25] and in-hand manipulation [23]. Furthermore, environmental constraints can also facilitate the learning of manipulation skills [26, 27, 28] for improved robustness and generalizability. In this work, we exploit environmental constraints to reduce the uncertainties caused by inaccurate sensing and approximate dynamics modeling, for robust compliant manipulation.
@@ III-A Preliminaries
We aim to find a sequence of Cartesian impedance controls, represented by a sequence of reference equilibrium poses $\left\{\bm{\xi}_{0},\cdots,\bm{\xi}_{T-1}\right\}$ of task frame, to move the robot gripper to reach the desired (yet initially occluded) grasp pose $\bm{X}^{g}\cdot\bm{X}^{*}$ from a starting collision-free pose $\bm{X}_{0}\in SE(3)$ through safe contacts and interactions with the environment.
III-B Exploring in Manipulation Funnel by Repetitions
The impedance controls are generated in an open loop and may not lead to a successful task completion during real executions, due to physical and perception uncertainties. However, as the failed execution does not greatly change the environment configuration, a repeated execution with some local randomness is still likely to achieve the task. To this end, our framework allows the robot to attempt more trials with repeated actions. Conceptually, as shown in Fig. 2, we can formulate a degenerate manipulation funnel that allows repeated actions to enter and navigate through. The entry of the funnel is defined by all initial task configurations (e.g., gripper poses, minor variations in object shape and poses); the exit of the funnel is all desired task completion states (e.g., stable grasps); the neck is the environmental or other task-implicit constraints that transition the manipulation towards the desired states. Each task repetition planned using environmental constraints is a series of motions that go inside the funnel from its entry towards the exit. Due to inaccurate modeling and uncertainties (as represented by the “flat” degenerate funnel neck in the figure), a repetition may not successfully find the funnel exit. However, the repetitions are always contained inside the funnel by not greatly disrupting the environment configuration and constraints. With more valid repetitions, the manipulation will finally locate the funnel exit to robustly finish the task.
@@ III-B Exploring in Manipulation Funnel by Repetitions
2:Task success ($true$ or $false$)
3:$k\leftarrow 0$triangleright Number of Repetitions
@@
6:    $\left\{\bm{X}^{1},\cdots,\bm{X}^{M}\right\}\leftarrow\textsc{Observe}(\hbox{})$triangleright Object Poses
@@
8:    $\left\{\widetilde{\bm{X}}_{t}\right\}_{t=1}^{T}\leftarrow\textsc{GeomPath}(\bm{X}_{0})$triangleright Sec. IV-B
9:    $\left\{\bm{\xi}_{t}\right\}_{t=0}^{T-1}\leftarrow\textsc{ImpdCtrl}(\left\{\widetilde{\bm{X}}_{t}\right\}_{t=1}^{T})$triangleright Sec. IV-C
@@
15:return $false$
Our proposed framework is outlined in Alg. 1. Specifically, the framework first observes the poses of objects in the environment. Provided with a desired grasp pose $\bm{X}^{*}$, a starting collision-free pre-grasp pose $\bm{X}_{0}$ of the gripper will be sampled around the workspace. To move the task frame of the robot gripper from $\bm{X}_{0}$ to the desired grasp $\bm{X}^{g}\cdot\bm{X}^{*}$, our collision-inclusive planner will then generate a sequence of desired impedance controls in two steps: 1) by exploiting environmental constraints (e.g., contacts with the objects or support surfaces), a geometric path will be planned for the task frame in Sec. IV-B; 2) Guided by the geometric path planned in the previous step, a sequence of impedance controls, represented by equilibrium poses for the task frame, will be generated via optimization in Sec. IV-C. The generated impedance controls will be commanded to the robot for execution. Meanwhile, if the task frame greatly deviates from the planned geometric path in Step 1), we will early terminate the execution by returning a $false$ in the $\textsc{Execute}(\cdot)$ function and skip to the next trial. After all impedance controls have been executed, the robot will close its gripper to grasp the target object. If the grasp is stable, as assessed by some off-the-shelf grasping criteria, the task is completed; otherwise, the robot will open the gripper, move to a newly sampled pre-grasp pose $\bm{X}_{0}$, and start the next trial. The repetition continues until the target object is stably grasped or the number of trials exceeds a preset limit $K_{\max}$.
@@ IV-A Geometric Representations of Robot and Objects
When planning and executing interactive robot motions, an efficient computation of the distance between the robot and environmental objects is needed. The exact computation of distance between complex mesh models in 3D space can be expensive and non-differentiable. Therefore, appropriate approximate geometric representations of the robot and objects are needed to facilitate efficient distance computation between shapes. As shown in Fig. 3 (left), we approximately represent the geometry of the robot gripper by a point cloud sampled on the gripper’s mesh model, i.e., $\mathcal{P}\subset\mathbb{R}^{3}$, where the coordinates of each point $\bm{p}\in\mathcal{P}$ is given in the task frame on the gripper. The points are sampled both on the surface and inside the volume of the gripper’s mesh model, i.e., $\mathcal{P}=\mathcal{P}^{s}\cup\mathcal{P}^{v}$ where $\mathcal{P}^{s}$ and $\mathcal{P}^{v}$ contain points sampled on the surface of and inside the mesh respectively.
Refer to captionTask Frame
Figure 3: Left: The gripper’s geometry is represented by a point cloud, randomly sampled on both the surface (blue) and inside the volume (red) of the mesh model. The task frame is defined at one fingertip. Right: The object (i.e., mustard bottle) is geometrically modeled by a signed distance field (SDF), where positive and negative values are visualized by yellow and green points respectively.
@@ IV-B Constraint-Exploiting Geometric Path Planning
Precisely placing the robot finger to a narrow gap for pre-grasp insertion merely based on visual perception is not reliable due to perceptual uncertainty. Instead, locally exploring and exploiting the environmental constraints (e.g., touching and sliding on the environmental surfaces) is more robust to such uncertainties for locating the gaps. For example, when inserting a finger into the gap between two tightly stacked books as shown in Fig. 4, directly hitting the finger toward the gap from in the air based on purely visual perception is likely to miss it since the gap’s location can be inaccurately observed. However, if we let the robot slide its finger on the book surface near the gap, the robot can more robustly find the gap by trapping its finger in the gap during sliding. Based on this intuition, we expect the task frame (at the fingertip) to follow a geometric path that actively exploits environmental constraints by contacting the object surfaces, as shown in Fig. 4. Meanwhile, we also want to reduce collisions between the environment and other parts (e.g., the gripper base) of the robot gripper, since unexpected contacts can easily deviate the gripper from its desired motion behavior while the robot moves compliantly. Taking into account both aspects, we propose to generate a desired geometric path for the gripper’s task frame in two steps:
@@ IV-B1 Positional Path-Finding via $A^{*}$ Algorithm
We solve the above path-finding problem by searching via an $A^{*}$ algorithm over the 3D workspace, as illustrated in Fig. 4 (right). Specifically, we discretize the workspace containing objects into a 3D grid with a resolution of $\Delta$. For the grids whose coordinates penetrate the environment by a value greater than $\delta_{g}$, i.e., violating the constraint in Eq. (5b), we regard this grid as being an obstacle to prevent the path from passing through it. For two neighbor grids with coordinates $\bm{p}$ and $\bm{p}^{\prime}$, the cost is defined as $c\left(\bm{p},\bm{p}^{\prime}\right)=w\cdot\phi\left(\bm{p}^{\prime}\right)^{2}+\lVert\bm{p}-\bm{p}^{\prime}\rVert^{2}$; and the heuristic function is simply the smoothing term (i.e., the path length), i.e., $h\left(\bm{p},\bm{p}^{\prime}\right)=\lVert\bm{p}-\bm{p}^{\prime}\rVert^{2}$, which underestimates the cost to ensure optimality of the found path. If the $A^{*}$ algorithm finds a path, represented by a sequence of traversed grids, we will interpolate it with $T$ waypoints to extract a positional path $\left\{\bm{p}_{1},\cdots,\bm{p}_{T}\right\}$ as the solution to Eq. (5).
Refer to caption9410
Figure 4: Positional path-finding via $A^{*}$ algorithm illustrated with a book-grasping example. Left: The fingertip of the gripper (i.e., the task frame) needs to follow a desired positional path (green), to exploit environmental constraints (i.e., contacts with object surfaces) for a more robust insertion into the gap between books. Right: The desired positional path (green) is planned on the 3D discretized workspace using an $A^{*}$ algorithm. The red grids penetrate the environment by a preset threshold $\delta_{g}$ and are regarded as obstacles. The colors of the blue grids (i.e., non-obstacle grids) represent their squared signed distance values, i.e., $\phi\left(\bm{p}\right)^{2}$, in $cm^{2}$.
@@
3:while $true$ do
4:    $path\leftarrow\textsc{AStarAlg}(\delta_{g})$triangleright Over a 3D Grid Workpsace
5:    if $path$ is not $null$ then

### v2

@@ Abstract
Robotic manipulation research has investigated contact-rich problems and strategies that require robots to intentionally collide with their environment, to accomplish tasks that cannot be handled by traditional collision-free solutions. By enabling compliant robot motions, collisions between the robot and its environment become more tolerable and can thus be exploited, but more physical uncertainties are introduced. To address contact-rich problems such as occluded object grasping while handling the involved uncertainties, we propose a collision-inclusive planning framework that can transition the robot to a desired task configuration via roughly modeled collisions absorbed by Cartesian impedance control. By strategically exploiting the environmental constraints and exploring inside a manipulation funnel formed by task repetitions, our framework can effectively reduce physical and perception uncertainties. With real-world evaluations on both single-arm and dual-arm setups, we show that our framework is able to efficiently address various realistic occluded grasping problems where a feasible grasp does not initially exist.
Index Terms:
Manipulation Planning, Grasping
@@ I Introduction
Different from traditional collision-free solutions where contacts between robots and the environment are prohibited [1, 2], recent research in robotic manipulation has investigated how robots can make effective collisions with the environment to reach desired task configurations in various contact-rich problems [3]. For a grasping example shown in Fig. 1, due to occlusions by surrounding objects, a feasible grasp configuration for the target object does not initially exist, and thus the robot has to collide with surrounding objects for a successful grasp. Like how humans grasp in such configurations, by allowing collisions and safe interactions between the gripper and the environment, the robot can insert its fingers into narrow gaps between objects to create free space for a desired grasp.
@@
Figure 1: Illustrations of occluded grasping tasks solved by our collision-inclusive planner. Top: A book is lying flat on a table surface, needing the robot gripper to scoop-grasp it by contacting the table. Bottom: A target book is tightly surrounded by other books on the shelf, requiring the robot gripper to insert its fingers into the gaps between books for a stable grasp.
@@
Motivated by the insights above, we propose a framework for occluded object grasping, for which all feasible grasps are initially blocked by the environment. Given mesh models of the robot gripper and objects, our framework is able to:
@@
    move the robot to its desired task configuration through trajectory generated by collision-inclusive planning;
@@
    generate effective compliant motions that exploit environmental constraints to reduce task uncertainties against inaccurate system modeling and perception;
@@ II Related Work
Contact-rich Manipulation. Robotic manipulation that requires contact with the environment has been extensively studied for different tasks and applications, such as wiping or polishing [10], peg-in-hole [11, 12], assembly [13], etc. Based on different task requirements and sensing capabilities, the strategies for planning and controlling robot motions need to incorporate different functionalities, such as estimation of contact state, control of contact forces, tolerance to inaccurate modeling, etc. Our framework uses a highly approximate system model to plan compliant motions, where controlling contact force is unnecessary for the tasks of our interest.
Occluded Object Grasping. When all feasible grasp configurations are occluded and traditional collision-free frameworks cannot solve the problem, solutions that actively leverage contacts to reconfigure objects are needed. Nonprehensile pre-grasp strategies such as pushing-based rearrangement [14, 15], sliding [16, 17, 18] and flipping [19, 20, 21] are developed via manually designed primitives or learned policies. In general, planning-based approaches [22] are more generalizable but rely on explicit contact modeling, making them sensitive to real-world uncertainties; while learning-based methods [23, 24] are more robust to such uncertainties but typically require large amounts of data to generalize. Unlike these existing methods which require significant reconfiguration of objects, our framework is effective in confined spaces without needing much object motions.
Exploitation of Environmental Constraints. Contacts with the environment can be leveraged as environmental constraints to help mitigate uncertainties [25]. This concept has been successfully applied in various manipulation problems to generate robust solutions in grasping [8, 26] and in-hand manipulation [27]. Furthermore, environmental constraints can also facilitate the learning of manipulation skills [28, 29, 30] for improved robustness and generalizability. In this work, we exploit environmental constraints in the form of a trajectory of contacts, to reduce the uncertainties caused by inaccurate sensing and approximate dynamics modeling.
@@ III-A Preliminaries
We aim to find a sequence of Cartesian impedance controls, defined by reference equilibrium poses $\left\{\bm{\xi}_{0},\cdots,\bm{\xi}_{T-1}\right\}$ of task frame, to move the robot gripper to reach the desired (yet initially occluded) grasp pose $\bm{X}^{g}\cdot\bm{X}^{*}$ from a starting collision-free pose $\bm{X}_{0}\in SE(3)$ through safe contacts and interactions with the environment. Unlike methods (e.g.,  [22, 23, 24]) that actively change object poses, our framework makes minimal changes to object configurations (i.e., $\bm{X}_{i}$), which is particularly effective when object motions are highly constrained (e.g., in a confined space with tightly packed objects, or when object motions can be blocked by another robot arm or environmental constraints).
@@
2:Task success (truetrue or falsefalse)
[…]
@@
6:  $\left\{\bm{X}^{1},\cdots,\bm{X}^{M}\right\}\leftarrow\textsc{Observe}(\hbox{})$ $\triangleright$ Object Poses
@@
8:  $\left\{\widetilde{\bm{X}}_{t}\right\}_{t=1}^{T}\leftarrow\textsc{GeomPath}(\bm{X}_{0},\bm{X}^{g}\cdot\bm{X}^{*})$ $\triangleright$ Sec. IV-B
9:  $\left\{\bm{\xi}_{t}\right\}_{t=0}^{T-1}\leftarrow\textsc{ImpdCtrl}(\left\{\widetilde{\bm{X}}_{t}\right\}_{t=1}^{T})$ $\triangleright$ Sec. IV-C
@@
15:return falsefalse
III-B Exploring in Manipulation Funnel by Repetitions
The impedance controls are generated in an open loop and may not lead to successful task completion during real executions, due to physical and perception uncertainties. However, as failed execution does not greatly change the environment configuration, a repeated execution with some local randomness is still likely to achieve the task. To this end, our framework allows the robot to attempt more trials with repeated actions. Conceptually, as shown in Fig. 2, we can formulate a degenerate manipulation funnel that allows repeated actions to enter and navigate through. The entry of the funnel is defined by all initial task configurations (e.g., gripper poses, minor variations in object shape and poses); the exit of the funnel is all desired task completion states (e.g., stable grasps); the neck is the environmental or other task-implicit constraints that transition the manipulation towards the desired states. Each task repetition planned using environmental constraints is a series of motions that go inside the funnel from its entry towards the exit. Due to inaccurate modeling and uncertainties (as represented by the “flat” degenerate funnel neck in the figure), a repetition may not successfully find the funnel exit. However, the repetitions are always contained inside the funnel by not greatly disrupting the environment configuration and constraints. With more valid repetitions, the manipulation will finally locate the funnel exit to robustly finish the task.
Our proposed framework is outlined in Alg. 1. Specifically, the framework first observes the poses of objects in the environment. Provided with a desired grasp pose $\bm{X}^{*}$, a starting collision-free pre-grasp pose $\bm{X}_{0}$ of the gripper will be sampled around the workspace. To move the task frame of the robot gripper from $\bm{X}_{0}$ to the desired grasp $\bm{X}^{g}\cdot\bm{X}^{*}$, our collision-inclusive planner will then generate a sequence of desired impedance controls in two steps: 1) by exploiting environmental constraints (e.g., contacts with the objects or support surfaces), a geometric path will be planned for the task frame in Sec. IV-B; 2) Guided by the geometric path planned in the previous step, a sequence of impedance controls, represented by equilibrium poses for the task frame, will be generated via optimization in Sec. IV-C. The generated impedance controls will be commanded to the robot for execution. Meanwhile, if the task frame greatly deviates from the planned geometric path in Step 1), we will early terminate the execution by returning a falsefalse in the $\textsc{Execute}(\cdot)$ function and skip to the next trial. After all impedance controls have been executed, the robot will close its gripper to grasp the target object. If the grasp is stable, as assessed by some off-the-shelf grasping criteria, the task is completed; otherwise, the robot will open the gripper, move to a newly sampled pre-grasp pose $\bm{X}_{0}$, and start the next trial. The repetition continues until the target object is stably grasped or the number of trials exceeds a preset limit $K_{\max}$.
@@ IV Collision-inclusive Planning
Refer to captionTask Frame
Figure 3: Left: The gripper’s geometry is represented by a point cloud: $1000$ points sampled on the surface (blue) and $50$ inside the volume (red) of the mesh model. The task frame is defined at one fingertip. Right: The object (i.e., mustard bottle) is modeled by a signed distance field (SDF), where positive and negative values are visualized by yellow and green points respectively.
@@ IV-A Geometric Representations of Robot and Objects
Approximate geometric representations are needed for efficient and differentiable distance computation between the robot and environment. As shown in Fig. 3 (left), we approximately represent the geometry of the robot gripper by a point cloud sampled on the gripper’s mesh model, i.e., $\mathcal{P}\subset\mathbb{R}^{3}$, where the coordinates of each point $\bm{p}\in\mathcal{P}$ is given in the task frame on the gripper. The points are sampled both on the surface and inside the volume of the gripper’s mesh model, i.e., $\mathcal{P}=\mathcal{P}^{s}\cup\mathcal{P}^{v}$. The surface points $\mathcal{P}^{s}$ should be sufficiently dense to capture the robot’s contact geometry, while the internal points $\mathcal{P}^{v}$ are kept sparse to reduce the computational cost of collision optimization.
@@ IV-B Constraint-Exploiting Geometric Path Planning
Refer to caption9410
Figure 4: Positional path-finding via $A^{*}$ algorithm illustrated with a book-grasping example. Left: The fingertip of the gripper (i.e., the task frame) needs to follow a desired positional path (green), to exploit environmental constraints (i.e., contacts with object surfaces) for a more robust insertion into the gap between books. Right: The desired positional path (green) is planned on the 3D discretized workspace using an $A^{*}$ algorithm. The red voxels penetrate the environment by a preset threshold $\delta_{g}$ and are regarded as obstacles. The colors of the blue voxels (i.e., non-obstacle voxels) represent their squared signed distance values, i.e., $\phi\left(\bm{p}\right)^{2}$, in $cm^{2}$.
Precisely placing the robot finger to a narrow gap for pre-grasp insertion merely based on visual perception is not reliable due to perceptual uncertainty. Instead, locally exploring and exploiting the environmental constraints (e.g., touching and sliding on the environmental surfaces) is more robust to such uncertainties for locating the gaps. For example, when inserting a finger into the gap between two tightly stacked books as shown in Fig. 4, directly hitting the finger toward the gap from in the air based on purely visual perception is likely to miss it since the gap’s location can be inaccurately observed. However, if we let the robot slide its finger on the book surface near the gap, the robot can more robustly find the gap by trapping its finger in the gap during sliding. Based on this intuition, we expect the task frame (at the fingertip) to follow a geometric path that actively exploits environmental constraints by contacting object surfaces, as shown in Fig. 4. Meanwhile, we also want to reduce collisions with other parts (e.g., the gripper base) of the robot gripper, since unexpected contacts can easily deviate the gripper from its desired motion behavior while the robot moves compliantly. Considering both aspects, we propose to generate a desired geometric path for the gripper’s task frame in two steps:
@@ IV-B1 Positional Path-Finding via $A^{*}$ Algorithm
We solve the above path-finding problem by searching via an $A^{*}$ algorithm over the 3D workspace, as illustrated in Fig. 4 (right). Specifically, we discretize the workspace containing objects into a 3D grid with resolution $\Delta$. For voxels whose coordinates penetrate the environment by a value greater than $\delta_{g}$, i.e., violating the constraint in Eq. (5b), we regard them as being an obstacle to prevent the path from passing through it. For two neighbor voxels with coordinates $\bm{p}$ and $\bm{p}^{\prime}$, the cost is defined as $c\left(\bm{p},\bm{p}^{\prime}\right)=w\cdot\phi\left(\bm{p}^{\prime}\right)^{2}+\lVert\bm{p}-\bm{p}^{\prime}\rVert^{2}$; and the heuristic function is simply the smoothing term (i.e., the path length), i.e., $h\left(\bm{p},\bm{p}^{\prime}\right)=\lVert\bm{p}-\bm{p}^{\prime}\rVert^{2}$, which underestimates the cost to ensure optimality of the found path. If a valid path (a sequence of traversed grid voxels) is found, we interpolate it with $T$ waypoints to extract a positional path $\left\{\bm{p}_{1},\cdots,\bm{p}_{T}\right\}$ as the solution to Eq. (5).
@@
3:while truetrue do
4:  $path\leftarrow\textsc{AStarAlg}(\delta_{g})$ $\triangleright$ Over a 3D Grid Workpsace
5:  if pathpath is not nullnull then